# A Comprehensive Survey of Website Fingerprinting Attacks and Defenses in Tor: Advances and Open Challenges

### v1

@@ V-A2 DL-based WF Attack models
RF [118] TAM and CNN DF [93] 95,000 96.77 40,000 >>85.00 NA
TinyWFP [118]
@@
WFCAT [115] $\times100$ 94.47 NA NA NA
CF [139]
@@
Wang14 [9]
AWF [94]
DF [93]
@@
AutoEncoder-based WF Attacks: In 2016, Abe et al. [10] first introduced DL into WF attacks, and proposed a fingerprinting attack based on the Stacked Denoising Autoencoder (SDAE) mechanism. The attack identifies websites accessed by users through analyzing traffic direction, even when the traffic is encrypted. This paper demonstrates the great potential of DL techniques in increasing the WF accuracy and achieves an 88% accuracy with the closed-world dataset. Later, Rimmer et al. [94] replaced the manual, expert-driven process of selecting traffic features with automatic feature learning via SDAE, CNN, and LSTM. The automated website fingerprinting (AWF) enables the attack to adapt better to traffic obfuscation techniques and changes over time. This approach also increases the practicality and stealth of WF attacks, especially in scenarios involving dynamic content or large-scale open-world environments. At the same time, Se Eun et al. [111] demonstrated that unsupervised DNNs, specifically autoencoders, can generate low-dimensional feature representations that improve the performance of the SOTA WF attacks. These AE-generated features reduce computational costs while maintaining or enhancing classifier accuracy.
@@
CNN-based WF Attacks: CNNs are originally designed for computer vision tasks, and have emerged as a powerful tool in WF attacks. Their ability to automatically extract hierarchical features from raw input makes them particularly effective at learning subtle patterns in encrypted or anonymized network traffic. In 2018, Sirinam et al. [93] introduced Deep Fingerprinting (DF), a CNN-based architecture designed to classify Tor traffic without requiring handcrafted features. DF achieves over 98% accuracy on their collected undefended Tor traffic, outperforming all prior WF attacks. In the same year, Bhat et al. [96] proposed Var-CNN, which combined CNNs with automated cumulative feature extraction (e.g., total packet counts, transmission time). Leveraging ResNet-18 and dilated causal convolutions, Var-CNN captured long-term dependencies in packet sequences, showing that combining packet direction and timing improves classification accuracy.
@@
Rahman et al. [12] later introduced burst-level timing features, which capture temporal relationships within and between bursts of traffic. This led to Tik-Tok, a WF attack using directional timing representations that combine packet timing and direction. To further exploit timing, Wang et al. [121] proposed a two-channel Temporal Convolutional Network (2ch-TCN) that separately models packet direction and timing sequences. On the Wang14 dataset [9], 2ch-TCN achieved 93.73% accuracy, outperforming CUMUL [11], Rimmer-CNN/SDAE [94], and Var-CNN [96]. Building on this, Shen et al. [152] developed BurNet, a fine-grained CNN-based WF method that processes unidirectional burst sequences rather than bidirectional packet sequences. BurNet achieved a precision and recall of 99.00%, making it effective in both local and remote attack settings. While defenses such as packet padding, traffic regularization, and delaying techniques reduce WF attack accuracy, Shen et al. [118] proposed Robust Fingerprinting (RF), which maintains strong performance under defenses. RF introduces a Traffic Aggregation Matrix (TAM) representation that encodes both packet direction and timing, allowing a CNN classifier to learn defense-resilient fingerprints. RF achieved an average accuracy improvement of 8.9% over Tik-Tok [93] against existing defenses.
@@
To address the computational cost of DL–based WF attacks, Tian et al. [130] introduced Tiny WFP, a lightweight method that applies wavelet-based dimensionality reduction and depthwise separable CNNs. By retaining only low-frequency wavelet coefficients, Tiny WFP achieved 99.2% accuracy in closed-world settings while being 81× smaller and 79× more efficient than DF. In parallel, Pan et al. [122] tackled the problem of limited training data. They combined CNNs with transfer learning to stabilize performance when traffic statistics vary irregularly over time. Their model integrated packet direction with timing sequences, significantly improving accuracy in few-shot scenarios. Gong et al. [115] argued that existing attacks underutilize timing features, which remain a persistent leakage source. They proposed WFCAT, which introduces an Inter-Arrival Time (IAT) histogram representation and CNN architecture with novel timing-aware blocks. WFCAT achieved over 59% accuracy against Surakav, outperforming RF and Tik-Tok by 28% and 48%, respectively, in the closed-world setting.
@@
Despite DL-based WF attacks achieving high accuracy, existing attacks suffer from high false positive rates, high computational cost, and low sensitivity. Chawla et al. [41] proposed ESPRESSO to improve the accuracy and efficiency of flow correlation attacks while reducing false positive rates and computational complexity compared to existing approaches. ESPRESSO utilizes an aggregated feature representation and employs Transformers for global processing to capture long-range dependencies. This method significantly outperforms DeepCoFFEA [153], especially in terms of reducing false positives and improving correlation accuracy.
@@
More recently, Zou et al. [40] introduced Relation-CNN, which enhances ResNet18 through three optimizations to preserve data fidelity and a feature-attention block to weight channel contributions. Relation-CNN improved prediction accuracy by up to 5% and 8% on defended and undefended datasets, respectively. To address the significant reduction in classification accuracy from trace adversarial examples, Shi [39] introduced a multiscale fingerprinting (MF). To effectively recognize the complex real patterns of anonymous traces under defenses, MF enhances feature extraction in two key ways. First, it slices each trace using multiscale time slots to construct multiscale trace matrices as inputs, thereby enriching temporal information. Second, it utilizes a sophisticated CNN structure with multiscale convolution layers to extract complicated features from changed patterns of varying lengths. Deng et al. [32] introduced Holmes, an early-stage WF attack that accurately identifies visited websites even when only a small portion of the page has loaded. Traditional DL-based WF attacks require complete page-loading traffic, making them unreliable under dynamic network conditions and WF defenses. Holmes overcomes this limitation by leveraging spatial-temporal distribution analysis to correlate early-stage traffic with pre-collected full traffic, significantly improving attack robustness.
@@
To address the concept drift problem, Xu et al. [42] introduce APWF, a parallel WF attack method that integrates an attention mechanism to address challenges like concept drift and small-sample inefficiency in traditional WF attacks. The APWF model combines CNN and BiGRU to capture spatial and temporal features of encrypted traffic, enhanced by a channel attention mechanism (CBAM) for improved feature learning. Additionally, the authors propose a fine-tuning method based on transfer learning to adapt to evolving website patterns with minimal labeled data. The method can maintain the accuracy at 92.4% in the scenario of 56 days between the training data and the target data, which is only 4% less loss compared to the instant attack.
@@
RNN-based WF Attacks: RNNs are designed to capture temporal dependencies in sequential data, making them suitable for modeling the time-series nature of network traffic, such as packet sequences in website visits. Compared to the outstanding performance of Autoencoders (AE) and CNN in the field of WF attacks, RNNs usually perform poorly and have higher time complexity  [94, 38]. However, RNNs are effective at capturing the temporal characteristics of website fingerprints and are often used in combination with CNNs. He et al. [151] introduced a complex deep neural network to automatically learn features and classify website fingerprints. The model uses a two-layer GRU network to extract the time feature and a 50-layer residual network to extract the spatial features of the website fingerprint. This method can achieve a classification accuracy rate of over 99%.
@@
TrafficSiam [136] AWF [94] NA 85.22 89.52 91.31 92.32 99.39 998.44 998.17 96.99
WF3A [138] Tik-Tok [12] 68.89 85.33 87.53 88.28 88.86 86.63 90.12 89.64 89.24
SMC [137]
Mixed ( [9]
 [94, 57])
@@
Sirinam et al. [95] first explored WF in Tor by proposing Triplet Fingerprinting (TF), a WF attack that leverages N-shot learning (NSL) to significantly reduce the amount of training data required. NSL allows the attacker to identify websites using only a few examples (e.g., 5 examples per website), making the attack more practical and portable. The base model refers to the CNN model that is used as the sub-network in the triplet networks. The K-NN model is implemented directly to measure the similarity between two different samples. This approach reduces the need for large datasets and frequent retraining, and achieves approximately 90% precision and 80% recall when tuned for precision. However, there is a large set of relevant auxiliary training samples for the model, which makes it unrealistic. To overcome this limitation, Chen et al. [97] improved Sirinam’s work by introducing a model-agnostic, efficient, and harmonious data augmentation (HDA) method, which aims to identify the websites a user visits by analyzing encrypted network traffic patterns, even in privacy-preserving networks like Tor. The HDA method offers three different data augmentation operations, including random rotation, masking, and mixing in intra-sample and inter-sample fashion. This improved attack model increases the accuracy from 87% to 90.60% under the dataset of Wang14 [9]. However, this method is unsatisfactory in small training data cases (e.g., less than 10 training samples per class). To overcome this limitation, Transfer Learning Fingerprinting Attack (TLFA) [133] was proposed by training a strong embedding model and leveraging classical machine learning models (e.g., SVM) on top. But this method remains inferior in the 1-shot case, especially for the open-world setting. Then, Chen et al. [127] introduce the understudied problem of few-shot WF attacks, emphasizing the need for adaptability to new websites with minimal labeled data. The paper addresses the challenge of few-shot WF attacks, where the goal is to identify visited websites from encrypted traffic data using only a few labeled training samples per website. This is a more realistic and scalable scenario compared to traditional WF attacks that assume abundant training data. The new few-shot model enhanced the accuracy up to 93% compared to previous works.
@@
Collecting large-scale and fresh traces is quite costly and unrealistic for DL-based WF attacks. When new traffic traces with different distributions or newly added monitored websites are introduced, DL-based WF attacks require a long bootstrap time for retraining. To address this limitation, Zou et al. [131] introduced a cross-domain, low-data WF attack called WFBDC (WF with Brownian Distance Covariance). WFBDC advances WF attacks by enabling efficient cross-domain few-shot learning with historical data. Its BDC-based approach enhances accuracy and reduces training time, demonstrating practicality for real-world deployment. Chen et al. [129] proposed a data augmentation technique that generates additional virtual traffic samples by extracting Trace-Forms (TForms), patterns derived from the structure of cell sequences in Tor traffic. By randomly refilling these TForms, they create new and meaningful samples that expand training data distribution. Through extensive experiments in both closed-world and open-world settings, TForm-RF is shown to significantly improve the classification accuracy of deep-learning-based WF attacks. Then, Xie et al. [139] introduced Contrastive Fingerprinting (CF), a WF attack leveraging contrastive learning and data augmentation to achieve high accuracy with minimal training traces. CF utilizes contrastive learning to produce a feature extractor in the pre-training stage. Then, the few-shot stage feeds the feature extractor with augmented traces and trains a classifier for testing users’ traces. Their work achieves a high accuracy of 90.4% in the closed-world scenario and distinguishes monitored websites with a high True Positive Rate of 91.2% in the open-world scenario. The previous few-shot-based WF attacks adopt complex metric strategies or perform time-consuming transfer learning, neither of which yields the most efficient performance in dynamic network environments. To address this issue, Lyu et al. [134] introduced Task-adaptive Feature Alignment Network (TFAN), a meta-learning-based approach for few-shot learning on WF attacks. TFAN regards the few-shot learning approach as a feature alignment problem in the class latent space, aiming to depict each location in the query feature map as a weighted sum of support features of a given class.
@@
Although existing few-shot WF attacks have shown some progress in addressing the data scarcity problem in WF attacks, they still have some limitations. Triplet Fingerprinting (TF) has a complex training process and does not fully leverage large auxiliary training data; TF and CWFA only focus on the feature space alignment between the originally and newly collected traces. TLFA and MBL methods use a fine-tuning process to obtain a task-specific classifier; however, they ignore the feature space alignment for the originally and newly collected traces. Zhou et al. [135] introduced a new few-shot WF attack method called Joint Alignment Networks (JAN) to address the limitations of existing methods in fine-grained feature alignment to improve the performance in few-shot WF attacks by jointly aligning the feature space and semantic space for originally and newly collected traces. This method enhances few-shot website fingerprinting (FS-WF) attacks, where an adversary aims to classify web traffic traces with minimal labeled data. Extensive experiments on public datasets show that JAN outperforms the state-of-the-art few-shot WF methods, especially in the difficult 1-shot tasks. Wang et al. [136] introduced TrafficSiam, a few-shot learning based WF attack that leverages contrastive learning and self-supervised pre-training to classify encrypted Tor traffic with minimal labeled data. Unlike traditional DL-based attacks [94, 93, 96, 95, 118] that require extensive labeled datasets, TrafficSiam pre-trains on unlabeled Tor traffic to learn robust feature representations, then fine-tunes with just a few labeled samples. This method achieves 92.32% accuracy in the closed-world setting. Wang et al. [138] introduce WF3A, a N-shot learning based WF attack that enhances attack accuracy in few-shot learning scenarios. Website fingerprinting is a privacy risk, allowing adversaries to identify websites visited by users over encrypted networks. Different from previous few-shot works, this work innovatively combines direction and length, designs and uses the fusion feature that covers more website identification information. To interpret the traffic patterns that can identify the website categories from the fusion feature, they introduce the Enhanced Channel Attention (ECA) and design a stronger feature extractor when constructing the WF model, which enhances the learning ability of the model by the channel attention mechanism.
@@
Traditional Few-Shot Website Fingerprinting (FSWF) attacks struggle with adaptability due to limited training data diversity [95, 138, 133, 131], leading to poor generalization across different network conditions. Tan et al. [137] introduced a joint learning framework to achieve the collusion FSWF attacks. The proposed method fuses the feature spaces of multiple user-side attackers to enhance the representation ability of the local model, and constructs a virtual fusion center to mitigate the impact of Non-IID (Non-Independent and Identically Distributed) data. Compared with the state-of-the-art method, the proposed method improves the accuracy by up to 13.02% in the closed-world setting and the AUC by up to 8.50% in the open-world setting, respectively. To address the challenge of WF in cross-environmental settings, where differences in browsers and proxy software cause feature drift and hinder model performance. Li et al. [140] proposed X-EPRINT, a systematic framework designed to generate cross-environmentally invariant features, apply potential-aware traffic resampling, and use inter-flow data augmentation for few-shot adaptation. The method tackles three core challenges: feature drift, the sampling dilemma, and few-shot generalization. Experimental evaluation across 18 scenarios shows that X-EPRINT achieves an F1-score of 0.719 in zero-shot recognition, 58.4% higher than the best baseline, and 0.925 in 3-shot adaptation, demonstrating strong robustness. However, its performance diminishes under severe feature drift, particularly with Firefox traffic, and the system still requires some resampling in new environments for optimal accuracy.
@@
Hybrid or other deep learning based WF Attacks: To enhance low-data training, Oh et al. [142] introduced Generative Adversarial Networks (GANs) for Data-Limited Fingerprinting (GANDaLF), which leverages GANs to generate synthetic data, enabling effective training of a deep neural network even with few real training samples. The paper evaluates GANDaLF in low-data scenarios, including fingerprinting both website index pages and non-index subpages, and achieves closed-world accuracy of 87% with just 20 instances per site (and 100 sites) in standard WF settings. In particular, GANDaLF outperforms TF by a 29% margin and Var-CNN by 38% for training sets using 20 instances per site. Guo et al. [87] introduced a homology analysis-based WF attack model, relying on a Convolutional Neural Network-Bidirectional Long Short-Term Memory (CNN-BiLSTM), targeting SSH and Shadowsocks encrypted traffic in anonymized networks. The method innovatively employs one-shot matrices to represent traffic features, enhancing fault tolerance and feature dimensionality, and achieves 94.8% and 98.1% accuracy in classifying SSH and Shadowsocks anonymous encrypted traffic.
@@
In 2022, Chen et al. [128] found that the current deep learning model is data-hungry when simulating the mapping relations of traffic and the website it belongs to, which may not be practical in reality. Then, they introduced the Send-and-Receive Pair (SRP) as a fundamental unit to analyze and generate bionic traffic traces for enhancing WF attacks on Tor. Their approach improves the Var-CNN model, achieving up to 50% higher accuracy in few-shot settings and outperforming transfer learning methods like TF in closed-world scenarios (95% accuracy). Wang et al. [124] analyzed the existing deep learning-based WF attack models and found that these attacks are sensitive to the specific information in the data and would result in less-than-ideal performance on data outside the training set or on data that is impacted by the effect of concept drift. They introduced snWF, which leverages an ensemble of deep learning models to improve the robustness and accuracy of identifying websites visited by users over encrypted connections. snWF introduces an ensemble technique called snapshot ensemble, where multiple snapshots of a deep neural network are taken during a single training run. These snapshots are then combined to reduce the variance of the model, making it more robust to concept drift and data outside the training set. This method achieves a TPR of 98.10% and a FPR of 5.70% with 400,000 open-word websites.

### v2

@@ V-A2 DL-based WF Attack models
RF [shen2023subverting] TAM and CNN DF [sirinam2018deep] 95,000 96.77 40,000 >>85.00 NA
TinyWFP [shen2023subverting]
@@
WFCAT [gong2024wfcat] $\times100$ 94.47 NA NA NA
CF [xie2024contrastive]
@@
Wang14 [wang2014effective]
AWF [rimmer2017automated]
DF [sirinam2018deep]
@@
AutoEncoder-based WF Attacks: In 2016, Abe et al. [abe2016fingerprinting] first introduced DL into WF attacks, and proposed a fingerprinting attack based on the Stacked Denoising Autoencoder (SDAE) mechanism. The attack identifies websites accessed by users through analyzing traffic direction, even when the traffic is encrypted. This paper demonstrates the great potential of DL techniques in increasing the WF accuracy and achieves an 88% accuracy with the closed-world dataset. Later, Rimmer et al. [rimmer2017automated] replaced the manual, expert-driven process of selecting traffic features with automatic feature learning via SDAE, CNN, and LSTM. The automated website fingerprinting (AWF) enables the attack to adapt better to traffic obfuscation techniques and changes over time. This approach also increases the practicality and stealth of WF attacks, especially in scenarios involving dynamic content or large-scale open-world environments. At the same time, Se Eun et al. [oh2017p] demonstrated that unsupervised DNNs, specifically autoencoders, can generate low-dimensional feature representations that improve the performance of the SOTA WF attacks. These AE-generated features reduce computational costs while maintaining or enhancing classifier accuracy.
@@
CNN-based WF Attacks: CNNs are originally designed for computer vision tasks, and have emerged as a powerful tool in WF attacks. Their ability to automatically extract hierarchical features from raw input makes them particularly effective at learning subtle patterns in encrypted or anonymized network traffic. In 2018, Sirinam et al. [sirinam2018deep] introduced Deep Fingerprinting (DF), a CNN-based architecture designed to classify Tor traffic without requiring handcrafted features. DF achieves over 98% accuracy on their collected undefended Tor traffic, outperforming all prior WF attacks. In the same year, Bhat et al. [bhat2018var] proposed Var-CNN, which combined CNNs with automated cumulative feature extraction (e.g., total packet counts, transmission time). Leveraging ResNet-18 and dilated causal convolutions, Var-CNN captured long-term dependencies in packet sequences, showing that combining packet direction and timing improves classification accuracy.
@@
Rahman et al. [rahman2019tik] later introduced burst-level timing features, which capture temporal relationships within and between bursts of traffic. This led to Tik-Tok, a WF attack using directional timing representations that combine packet timing and direction. To further exploit timing, Wang et al. [wang20202ch] proposed a two-channel Temporal Convolutional Network (2ch-TCN) that separately models packet direction and timing sequences. On the Wang14 dataset [wang2014effective], 2ch-TCN achieved 93.73% accuracy, outperforming CUMUL [panchenko2016website], Rimmer-CNN/SDAE [rimmer2017automated], and Var-CNN [bhat2018var]. Building on this, Shen et al. [shen2021efficient] developed BurNet, a fine-grained CNN-based WF method that processes unidirectional burst sequences rather than bidirectional packet sequences. BurNet achieved a precision and recall of 99.00%, making it effective in both local and remote attack settings. While defenses such as packet padding, traffic regularization, and delaying techniques reduce WF attack accuracy, Shen et al. [shen2023subverting] proposed Robust Fingerprinting (RF), which maintains strong performance under defenses. RF introduces a Traffic Aggregation Matrix (TAM) representation that encodes both packet direction and timing, allowing a CNN classifier to learn defense-resilient fingerprints. RF achieved an average accuracy improvement of 8.9% over Tik-Tok [sirinam2018deep] against existing defenses.
@@
To address the computational cost of DL–based WF attacks, Tian et al. [tian2023tiny] introduced Tiny WFP, a lightweight method that applies wavelet-based dimensionality reduction and depthwise separable CNNs. By retaining only low-frequency wavelet coefficients, Tiny WFP achieved 99.2% accuracy in closed-world settings while being 81× smaller and 79× more efficient than DF. In parallel, Pan et al. [pan2023practical] tackled the problem of limited training data. They combined CNNs with transfer learning to stabilize performance when traffic statistics vary irregularly over time. Their model integrated packet direction with timing sequences, significantly improving accuracy in few-shot scenarios. Gong et al. [gong2024wfcat] argued that existing attacks underutilize timing features, which remain a persistent leakage source. They proposed WFCAT, which introduces an Inter-Arrival Time (IAT) histogram representation and CNN architecture with novel timing-aware blocks. WFCAT achieved over 59% accuracy against Surakav, outperforming RF and Tik-Tok by 28% and 48%, respectively, in the closed-world setting.
@@
Despite DL-based WF attacks achieving high accuracy, existing attacks suffer from high false positive rates, high computational cost, and low sensitivity. Chawla et al. [chawla2024espresso] proposed ESPRESSO to improve the accuracy and efficiency of flow correlation attacks while reducing false positive rates and computational complexity compared to existing approaches. ESPRESSO utilizes an aggregated feature representation and employs Transformers for global processing to capture long-range dependencies. This method significantly outperforms DeepCoFFEA [oh2022deepcoffea], especially in terms of reducing false positives and improving correlation accuracy.
@@
More recently, Zou et al. [zou2024relation] introduced Relation-CNN, which enhances ResNet18 through three optimizations to preserve data fidelity and a feature-attention block to weight channel contributions. Relation-CNN improved prediction accuracy by up to 5% and 8% on defended and undefended datasets, respectively. To address the significant reduction in classification accuracy from trace adversarial examples, Shi [shi2025multiscale] introduced a multiscale fingerprinting (MF). To effectively recognize the complex real patterns of anonymous traces under defenses, MF enhances feature extraction in two key ways. First, it slices each trace using multiscale time slots to construct multiscale trace matrices as inputs, thereby enriching temporal information. Second, it utilizes a sophisticated CNN structure with multiscale convolution layers to extract complicated features from changed patterns of varying lengths. Deng et al. [deng2024robust] introduced Holmes, an early-stage WF attack that accurately identifies visited websites even when only a small portion of the page has loaded. Traditional DL-based WF attacks require complete page-loading traffic, making them unreliable under dynamic network conditions and WF defenses. Holmes overcomes this limitation by leveraging spatial-temporal distribution analysis to correlate early-stage traffic with pre-collected full traffic, significantly improving attack robustness.
@@
To address the concept drift problem, Xu et al. [xu2025apwf] introduce APWF, a parallel WF attack method that integrates an attention mechanism to address challenges like concept drift and small-sample inefficiency in traditional WF attacks. The APWF model combines CNN and BiGRU to capture spatial and temporal features of encrypted traffic, enhanced by a channel attention mechanism (CBAM) for improved feature learning. Additionally, the authors propose a fine-tuning method based on transfer learning to adapt to evolving website patterns with minimal labeled data. The method can maintain the accuracy at 92.4% in the scenario of 56 days between the training data and the target data, which is only 4% less loss compared to the instant attack.
@@
RNN-based WF Attacks: RNNs are designed to capture temporal dependencies in sequential data, making them suitable for modeling the time-series nature of network traffic, such as packet sequences in website visits. Compared to the outstanding performance of Autoencoders (AE) and CNN in the field of WF attacks, RNNs usually perform poorly and have higher time complexity  [rimmer2017automated, hong2024website]. However, RNNs are effective at capturing the temporal characteristics of website fingerprints and are often used in combination with CNNs. He et al. [he2018deep] introduced a complex deep neural network to automatically learn features and classify website fingerprints. The model uses a two-layer GRU network to extract the time feature and a 50-layer residual network to extract the spatial features of the website fingerprint. This method can achieve a classification accuracy rate of over 99%.
@@
TrafficSiam [wang2024trafficsiam] AWF [rimmer2017automated] NA 85.22 89.52 91.31 92.32 99.39 998.44 998.17 96.99
WF3A [wang2024wf3a] Tik-Tok [rahman2019tik] 68.89 85.33 87.53 88.28 88.86 86.63 90.12 89.64 89.24
SMC [tan2024adaptability]
Mixed ( [wang2014effective]
 [rimmer2017automated, gong2020zero])
@@
Sirinam et al. [sirinam2019triplet] first explored WF in Tor by proposing Triplet Fingerprinting (TF), a WF attack that leverages N-shot learning (NSL) to significantly reduce the amount of training data required. NSL allows the attacker to identify websites using only a few examples (e.g., 5 examples per website), making the attack more practical and portable. The base model refers to the CNN model that is used as the sub-network in the triplet networks. The K-NN model is implemented directly to measure the similarity between two different samples. This approach reduces the need for large datasets and frequent retraining, and achieves approximately 90% precision and 80% recall when tuned for precision. However, there is a large set of relevant auxiliary training samples for the model, which makes it unrealistic. To overcome this limitation, Chen et al. [chen2021few] improved Sirinam’s work by introducing a model-agnostic, efficient, and harmonious data augmentation (HDA) method, which aims to identify the websites a user visits by analyzing encrypted network traffic patterns, even in privacy-preserving networks like Tor. The HDA method offers three different data augmentation operations, including random rotation, masking, and mixing in intra-sample and inter-sample fashion. This improved attack model increases the accuracy from 87% to 90.60% under the dataset of Wang14 [wang2014effective]. However, this method is unsatisfactory in small training data cases (e.g., less than 10 training samples per class). To overcome this limitation, Transfer Learning Fingerprinting Attack (TLFA) [chen2021few2] was proposed by training a strong embedding model and leveraging classical machine learning models (e.g., SVM) on top. But this method remains inferior in the 1-shot case, especially for the open-world setting. Then, Chen et al. [chen2022few] introduce the understudied problem of few-shot WF attacks, emphasizing the need for adaptability to new websites with minimal labeled data. The paper addresses the challenge of few-shot WF attacks, where the goal is to identify visited websites from encrypted traffic data using only a few labeled training samples per website. This is a more realistic and scalable scenario compared to traditional WF attacks that assume abundant training data. The new few-shot model enhanced the accuracy up to 93% compared to previous works.
@@
Collecting large-scale and fresh traces is quite costly and unrealistic for DL-based WF attacks. When new traffic traces with different distributions or newly added monitored websites are introduced, DL-based WF attacks require a long bootstrap time for retraining. To address this limitation, Zou et al. [zou2022efficient] introduced a cross-domain, low-data WF attack called WFBDC (WF with Brownian Distance Covariance). WFBDC advances WF attacks by enabling efficient cross-domain few-shot learning with historical data. Its BDC-based approach enhances accuracy and reduces training time, demonstrating practicality for real-world deployment. Chen et al. [chen2022tform] proposed a data augmentation technique that generates additional virtual traffic samples by extracting Trace-Forms (TForms), patterns derived from the structure of cell sequences in Tor traffic. By randomly refilling these TForms, they create new and meaningful samples that expand training data distribution. Through extensive experiments in both closed-world and open-world settings, TForm-RF is shown to significantly improve the classification accuracy of deep-learning-based WF attacks. Then, Xie et al. [xie2024contrastive] introduced Contrastive Fingerprinting (CF), a WF attack leveraging contrastive learning and data augmentation to achieve high accuracy with minimal training traces. CF utilizes contrastive learning to produce a feature extractor in the pre-training stage. Then, the few-shot stage feeds the feature extractor with augmented traces and trains a classifier for testing users’ traces. Their work achieves a high accuracy of 90.4% in the closed-world scenario and distinguishes monitored websites with a high True Positive Rate of 91.2% in the open-world scenario. The previous few-shot-based WF attacks adopt complex metric strategies or perform time-consuming transfer learning, neither of which yields the most efficient performance in dynamic network environments. To address this issue, Lyu et al. [lyu2024tfan] introduced Task-adaptive Feature Alignment Network (TFAN), a meta-learning-based approach for few-shot learning on WF attacks. TFAN regards the few-shot learning approach as a feature alignment problem in the class latent space, aiming to depict each location in the query feature map as a weighted sum of support features of a given class.
@@
Although existing few-shot WF attacks have shown some progress in addressing the data scarcity problem in WF attacks, they still have some limitations. Triplet Fingerprinting (TF) has a complex training process and does not fully leverage large auxiliary training data; TF and CWFA only focus on the feature space alignment between the originally and newly collected traces. TLFA and MBL methods use a fine-tuning process to obtain a task-specific classifier; however, they ignore the feature space alignment for the originally and newly collected traces. Zhou et al. [zhou2024joint] introduced a new few-shot WF attack method called Joint Alignment Networks (JAN) to address the limitations of existing methods in fine-grained feature alignment to improve the performance in few-shot WF attacks by jointly aligning the feature space and semantic space for originally and newly collected traces. This method enhances few-shot website fingerprinting (FS-WF) attacks, where an adversary aims to classify web traffic traces with minimal labeled data. Extensive experiments on public datasets show that JAN outperforms the state-of-the-art few-shot WF methods, especially in the difficult 1-shot tasks. Wang et al. [wang2024trafficsiam] introduced TrafficSiam, a few-shot learning based WF attack that leverages contrastive learning and self-supervised pre-training to classify encrypted Tor traffic with minimal labeled data. Unlike traditional DL-based attacks [rimmer2017automated, sirinam2018deep, bhat2018var, sirinam2019triplet, shen2023subverting] that require extensive labeled datasets, TrafficSiam pre-trains on unlabeled Tor traffic to learn robust feature representations, then fine-tunes with just a few labeled samples. This method achieves 92.32% accuracy in the closed-world setting. Wang et al. [wang2024wf3a] introduce WF3A, a N-shot learning based WF attack that enhances attack accuracy in few-shot learning scenarios. Website fingerprinting is a privacy risk, allowing adversaries to identify websites visited by users over encrypted networks. Different from previous few-shot works, this work innovatively combines direction and length, designs and uses the fusion feature that covers more website identification information. To interpret the traffic patterns that can identify the website categories from the fusion feature, they introduce the Enhanced Channel Attention (ECA) and design a stronger feature extractor when constructing the WF model, which enhances the learning ability of the model by the channel attention mechanism.
@@
Traditional Few-Shot Website Fingerprinting (FSWF) attacks struggle with adaptability due to limited training data diversity [sirinam2019triplet, wang2024wf3a, chen2021few2, zou2022efficient], leading to poor generalization across different network conditions. Tan et al. [tan2024adaptability] introduced a joint learning framework to achieve the collusion FSWF attacks. The proposed method fuses the feature spaces of multiple user-side attackers to enhance the representation ability of the local model, and constructs a virtual fusion center to mitigate the impact of Non-IID (Non-Independent and Identically Distributed) data. Compared with the state-of-the-art method, the proposed method improves the accuracy by up to 13.02% in the closed-world setting and the AUC by up to 8.50% in the open-world setting, respectively. To address the challenge of WF in cross-environmental settings, where differences in browsers and proxy software cause feature drift and hinder model performance. Li et al. [li2025cross] proposed X-EPRINT, a systematic framework designed to generate cross-environmentally invariant features, apply potential-aware traffic resampling, and use inter-flow data augmentation for few-shot adaptation. The method tackles three core challenges: feature drift, the sampling dilemma, and few-shot generalization. Experimental evaluation across 18 scenarios shows that X-EPRINT achieves an F1-score of 0.719 in zero-shot recognition, 58.4% higher than the best baseline, and 0.925 in 3-shot adaptation, demonstrating strong robustness. However, its performance diminishes under severe feature drift, particularly with Firefox traffic, and the system still requires some resampling in new environments for optimal accuracy.
@@
Hybrid or other deep learning based WF Attacks: To enhance low-data training, Oh et al. [oh2021gandalf] introduced Generative Adversarial Networks (GANs) for Data-Limited Fingerprinting (GANDaLF), which leverages GANs to generate synthetic data, enabling effective training of a deep neural network even with few real training samples. The paper evaluates GANDaLF in low-data scenarios, including fingerprinting both website index pages and non-index subpages, and achieves closed-world accuracy of 87% with just 20 instances per site (and 100 sites) in standard WF settings. In particular, GANDaLF outperforms TF by a 29% margin and Var-CNN by 38% for training sets using 20 instances per site. Guo et al. [guo2021website] introduced a homology analysis-based WF attack model, relying on a Convolutional Neural Network-Bidirectional Long Short-Term Memory (CNN-BiLSTM), targeting SSH and Shadowsocks encrypted traffic in anonymized networks. The method innovatively employs one-shot matrices to represent traffic features, enhancing fault tolerance and feature dimensionality, and achieves 94.8% and 98.1% accuracy in classifying SSH and Shadowsocks anonymous encrypted traffic.
@@
In 2022, Chen et al. [chen2022srp] found that the current deep learning model is data-hungry when simulating the mapping relations of traffic and the website it belongs to, which may not be practical in reality. Then, they introduced the Send-and-Receive Pair (SRP) as a fundamental unit to analyze and generate bionic traffic traces for enhancing WF attacks on Tor. Their approach improves the Var-CNN model, achieving up to 50% higher accuracy in few-shot settings and outperforming transfer learning methods like TF in closed-world scenarios (95% accuracy). Wang et al. [wang2022snwf] analyzed the existing deep learning-based WF attack models and found that these attacks are sensitive to the specific information in the data and would result in less-than-ideal performance on data outside the training set or on data that is impacted by the effect of concept drift. They introduced snWF, which leverages an ensemble of deep learning models to improve the robustness and accuracy of identifying websites visited by users over encrypted connections. snWF introduces an ensemble technique called snapshot ensemble, where multiple snapshots of a deep neural network are taken during a single training run. These snapshots are then combined to reduce the variance of the model, making it more robust to concept drift and data outside the training set. This method achieves a TPR of 98.10% and a FPR of 5.70% with 400,000 open-word websites.